The angle between the vectors (i^+j^+k^) and (i^-j^-k^) is 
  1. sin-1√8/3
  2. Sin-1(1/3)
  3. Cos-1√8/3
  4. Cos-1√(8/3)

Grade:12th pass

1 Answers

Harshit Singh
3 years ago


Dear Student

cos x = Dot product of vector A and B/ product of their magnitudes
dot product of A and B= 1-1-1= -1
magnitude of vector A =magnitude of vector B= √3

cos x = (-1)/(√3.√3)
= -1/3

sin^2 x = 1 – cos^2 x
sin x = √8/3
x = sin inverse (√8/3)

Thanks